OpenStudy (anonymous):

Find the exact area under one arch of the curve y(t) = ksin(wt) where k and w are positive constants. Note: Use lower case k and w.

OpenStudy (anonymous):

im stuck

zepdrix (zepdrix):

Hmm let's start by finding the correct boundaries for one arch.

zepdrix (zepdrix):

\[\Large\bf\sf y(t)\quad=\quad k \sin(\omega t)\]The period is given by:\[\Large\bf\sf P=\frac{2\pi}{\omega}\]

zepdrix (zepdrix):

Which consists of BOTH arches, right? So we want half of that.

zepdrix (zepdrix):

Sine goes through the origin, ok good we can start at 0. So our boundaries will be from \(\Large\bf\sf0\) to \(\Large\bf\sf \dfrac{\pi}{\omega}\)

zepdrix (zepdrix):

Any confusion there? :o

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits_0^{\pi/\omega}k \sin(\omega t)\;dt\]

zepdrix (zepdrix):

Bring the k outside, it's just a constant.\[\Large\bf\sf k\int\limits\limits_0^{\pi/\omega} \sin(\omega t)\;dt\]From there you can apply a u-substitution if the omega factor on t is going to confuse you.

zepdrix (zepdrix):

Oh if it's supposed to be a w, then my bad.

OpenStudy (anonymous):

thank you sorry I was working it out on my own and I managed to get it without your help my bad i am so used to people ignoring my questions sorry